OpenStudy (anonymous):

Find cos θ if sin θ = -5/13 and tan θ > 0.

myininaya (myininaya):

If sine is negative and tangent is positive, then cosine is ____?____ tangent=sine/cosine=negative/cosine=positive =>cosine is ___?____

OpenStudy (anonymous):

positive?

myininaya (myininaya):

So negative/positive is positive No...

OpenStudy (anonymous):

no its negative

myininaya (myininaya):

Right. You want tangent to be positive not negative

OpenStudy (anonymous):

positive

myininaya (myininaya):

So you need cosine to be negative in order for tangent to be positive since sine is negative.

myininaya (myininaya):

So we are in the (-,-) quadrant.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

correct

myininaya (myininaya):

 \[\sin( \theta)=-5/13 \text{ is given to you} \]

OpenStudy (anonymous):

so then use pythagorean theorem to fine cosine

myininaya (myininaya):

Find the other side using the Pythagorean thm.

OpenStudy (anonymous):

which would end up being -12/13

myininaya (myininaya):

the other side is 12 is what you are saying and no i thought you were looking for cosine...

myininaya (myininaya):

oops I labeled the triangle wrong a little bit.

OpenStudy (anonymous):

i am

myininaya (myininaya):